What is the final velocity of a ball rolling down an inclined plane with a velocity of 4.0m/s and accelerate at 2.0m/s² in 5.0s?


1
Expert's answer
2021-07-28T09:11:03-0400

By definition, the acceleration is given as follows:


"a = \dfrac{v_f - v_i}{t}"

where "v_f, \space v_i = 4.0m/s" are the final and initial velocities of the ball respectively, and "t = 5.0s" is the time. Expressing "v_f" and substituting "a = 2.0m\/s^2", obtain:


"v_f = v_i + at\\
v_f = 4.0m/s + 2.0m/s^2\cdot 5.0s =14m/s"

Answer. 14 m/s.